Okay here is the one that I just cant seem to find a starting point. If 5 times the lesser of two consecutive integers is added to three times the greater, the result is 59. Find the integers. I have 5x-2x=3times =59 and I believe I kind have drove a car over the cliff. I am really struggling with the word problems.
Answer by richwmiller(17219)

Okay here is the one that I just cant seem to find a starting point.
Let's call them n and n+1
If 5 times the lesser of two consecutive integers 5n
is added +
to three times the greater 3*(n+1)
the result is =
59 = 59.
5n+3*(n+1)=59
Find the integers. I have 5x-2x=3times =59 and I believe I kind have drove a car over the cliff. I am really struggling with the word problems.
Most students have trouble with word problems.
5n+3*(n+1)=59